35 L of oxygen at 85 C and 92,000 N m-2 is added to a 40 L stainless steel cylinder
containing carbon dioxide at 40 C and 70,000 N m-2. Calculate the final pressure in
the cylinder if after the addition the final temperature is 55 C.
1
Expert's answer
2013-12-03T11:20:31-0500
There is no chemical reactions in system, so we can calculate its pressure from physical laws.
Calculate the pressures of gases at the final conditions:

Oxygen:
P1V1/T1 = P2V2/T2
P2(O2) = P1V1T2/V2T1 = 92,000*35*328/(40*358) = 73,750 N/m^2

Carbon dioxide:
P1/T1 = P2/T2
P2(CO2) = P1T2/T1 = 73,350 N/m^2

The total pressure in systems is the sum of partial pressures of components:

P = P2(O2) + P2(CO2) = 147,200 N/m^2 = 147,000 N/m^2